OpenStudy (anonymous):

what math teacher made up this question? Glenn needs 20 minutes to clean his room. If his little sister Veronika pulls things out while he is cleaning, it takes 60 minutes for him to clean the room. How many minutes does it take for Veronika to pull everything out?

OpenStudy (anonymous):

lol where did you find that?

OpenStudy (anonymous):

I was asking my teacher for bonus questions and she told me try this

OpenStudy (anonymous):

i wrote that

OpenStudy (anonymous):

i mean i wrote "what math teacher made up this question" the rest i copied from someone here who actually asked it who is your math teacher?

OpenStudy (anonymous):

I had wrote that because this question is really dumb and i really want to know who made up this question

OpenStudy (anonymous):

someone with no common sense, that's who

OpenStudy (triciaal):

it's a bonus question gives the student something to do

OpenStudy (anonymous):

i says five second

OpenStudy (anonymous):

that was my answer but it's saying that's wrong

OpenStudy (triciaal):

i think 3 mins

OpenStudy (anonymous):

that is how long it would take me anyways the math teacher answer is the solution to \[\frac{1}{20}-\frac{1}{x}=\frac{1}{60}\] or \(x=30\)